CAT Chat Minutes

December 6, 2004

 

Information and Follow-up
Last week the question was asked about when roads that had been cut through to allow for CNM piping to be laid, were going to be repaired.   Bob Whitman sent e-mail today explaining the schedule for road repairs and is pasted here:

Subject: APS site road repairs

From: Bob Whitman <rrw@aps.anl.gov>

Date: Fri, 03 Dec 2004 12:03:17 -0600

Organization: Argonne National Laboratory -- Advanced Photon Source

On Monday, 12/6, weather permitting, the repair all of the road cuts that were made for the installation of the CNM's gas and electric supply lines will begin. This work will start at approximately 9:00AM and continue for approximately three days.

The areas that will be repaired are as follows:  Rock Road at Kearney, and Near building 450 and 401, Building 401 east parking lot, the roads to and from the Argonne Guest House, Building 302 parking lot, Building 301 driveway, sidewalk, and parking lot.

As the contractors move from one repair area to the next they will be closing roads, and controlling the traffic as necessary. An effort will be made to maintain traffic flow by controlling the traffic rather than closing the roads.

The south four spaces of each row in the southeast area of the 401 parking lot (adjacent to building 450) will be blocked off for the duration of the repair work and the road adjacent to building 450 will be closed for a short time.

Please observe all road closure and detour signs and flagmen at the various work sites. SAFETY FIRST AS ALWAYS AT APS AND ANL!!!!   Please inform all affected parties of this work.

While checking the APS web pages over the weekend, a user was unable to download the images.   What happened?

We are in the process of adjusting our firewalls; you should be able to download images now.

Request for detectors for winter runs by January 13, 2005.   See email regarding detector pool reservation from – Steve Ross-    Dean Haeffner asked why the deadline  is so close to the run – isn’t it almost too late?   He asks that APS please consider this for the next time, realizing that it is too late now to change this timetable.  

General Information
Machine Stats - R. Klaffky:   Run 3-2004 Operations Statistics
            Delivered beam:   1207 hrs.
            Percentage of scheduled time:   98.57%
            Downtime during period:   17.51 hrs.
            Mean time between faults:   121.37 hrs.

Faults:

# 18   11/26 08:00   To   11/30 15:59   103.99     Int Dump: End of Period     0.00
                                                                        0.67
# 19   12/02 08:40   To   12/03 08:38    23.96    34-ID BPLD Trip     [DIA]   0.39 X-Ray BPM pre-amp chassis error, refill, top-up
# 20   12/03 09:01   To   12/06 08:04    71.05    Fill Ongoing                0.00  

Beamline Downtime by APS Groups:   Summary by APS groups for period 11/26/2004 to 12/03/2004

DIN

Date

Beamline

Status

Duration

Group

Description

Resolution

488

11/29/04

02-BM

Open

30 min

AOD-IT

Beamline server is down, and users cannot save data.

Users reported that Dave Paterson was working on the problem from home. Rodney reached Dave Wallis; he reported that server is running again.

489

11/29/04

02-ID

Open

3 hr 0 min

AOD-IT

Beamline server is down, and users cannot save data. David Paterson reports that downtime is actually three hours, because an entire scan was lost.

Users reported that Dave Paterson was working on the problem from home. Rodney reached Dave Wallis; he reported that server is running again.

Injector Availability:   96.23%

Out of range at:   12/03/2004 14:56:565 to 12/03/2004 18:41:35 | 224.67 mins.

Booster ramp supply replaced with PFS performing LOTO

Cybersecurity – K. Sidorowicz:  There was a security break-in of 2BM website recently. APS-IT group was managing the website but the web content is managed by sector 2. A user notified Gabrielle Long, XFD, who then notified the IT group and the website was taken down. Argonne Cyber security was called by the FBI who was already investigating the incident.  The individual was able to drop in a PHP BOT program and run it and then upload a shell program. The IP number of the machine used for the break-in is from Indonesia.  We are now monitoring the beamline websites and putting together some security training sessions.

Safety – E. Chang:   Safety inspections scheduled for December are 431, 432, and 433. Bruce Glagola will also participate in inspections.  

We are missing an electrical tool tester (sensitive item number 009) which is due now for calibration.   Please notify Elroy if you know where this tool is located.

On December 9, 2004 in conference room B400 beginning at 2:00 p.m., the Laser Safety team will meet to discuss the charter that was submitted. If you have any comments, contact Elroy before this meeting.

The eight-hour course “Electrical Safety” training is not entirely applicable to all of our needs.   Nineteen people are listed as attending thus far.   Elroy Chang and Richard Hislop are working to provide to the Laboratory more issues to be addressed in a subset course.  

Maintenance Period:   Once again, we remind the groups that we have no overtime allowance for HP technicians, therefore do not request work done with radioactive samples on weekends.   Remember, too, to put in your requests for work during the shutdown to your floor coordinators as soon as possible.